8. Two boys are pulling a box across a horizontal floor. If F1-50 N,300 N of E and F2 100 N, 30° E of N, what is the resultant force on the box?
Answer #1


writing the components for forces


F1x = F1*cos30 = 50*cos30 = 43.3 N


F1y = F1*sin30 = 50*sin30 = 25 N

F2x = F2*sin30 = 100*sin30 = 50 N


F2y = F2*cos300 = 100*cos30 = 86.6 N


Rx = F1x + F2x = 93.3 N

Ry = F1y + F2y = 111.6 N

resultant force R = sqrt(Rx^2 +RY^2) = sqrt(93.3^2+111.6^2) = 145.46 N

direction


tan^-1(111.6/93.3) = 50.1 norht of east